# Online Multi-modal Root Cause Analysis

### v1

@@ 3.2 Long-term Temporal Causal Structure Learning
where $\bm{f}(t)\in\mathbb{R}^{K}$ represents the 1-D kernel, $d$ is the dilation factor controlling the skipping distance, $\odot$ denotes the Hadamard product, $v\in\{M,L\}$, $\sigma(x)=\frac{1}{(1+e^{-x})}$ is the sigmoid function, and $\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ is the tanh function. $\bm{f}_{1}$, $\bm{f}_{2}$, $\bm{f}_{3}$, and $\bm{f}_{4}$ are 1-D kernels of the dilated convolution networks. $\bm{H}_{v}^{0}\in\mathbb{R}^{n\times d_{v}\times T_{3}}$ and $\bm{H}_{v}^{i}\in\mathbb{R}^{n\times d_{v}\times T_{4}}$ represent the historical time series and the $i^{\textrm{th}}$ batch of streaming time series for the modality $v$, respectively. $T_{3}$ and $T_{4}$ are the output dimensions of the dilated convolution networks. Additionally, $MFL(\cdot)$ denotes the representation learning with multi-factor attention module, aiming to encode the correlation of different metrics into the representations $\hat{\bm{O}}_{v}^{0}\in\mathbb{R}^{n\times d_{v}\cdot T_{3}}$ and $\hat{\bm{O}}_{v}^{i}\in\mathbb{R}^{n\times d_{v}\cdot T_{4}}$, which will be introduced in the next subsection. By stacking dilated causal convolution layers, the model’s receptive field grows exponentially, allowing for longer time series with fewer layers and thereby reducing computation costs. We validate the efficiency of dilated convolutional operations in our experiments (see Subsection 4.2) by comparing their computational costs with those of VAR-based methods.
To learn the causal relationship among system entities, we aggregate information from neighbors via a graph neural network (Kipf & Welling, 2017) and mimic fault propagation through a message-passing mechanism:
$\tilde{\bm{X}}_{v}^{0}$ $=\sigma_{2}(\mathcal{A}_{old}(\hat{\bm{O}}_{v}^{0}\oplus\bm{N}^{0}_{v})\bm{W}^{1}),\text{where }\bm{N}^{0}_{v}[j]=\frac{1}{|\mathcal{N}_{j}|}\sum_{k\in N_{j}}\hat{\bm{O}}_{v}^{0}[k]$ (6)
$\tilde{\bm{X}}_{v}^{i}$ $=\sigma_{2}((\mathcal{A}_{old}+\Delta\mathcal{A}_{v})(\hat{\bm{O}}_{v}^{i}\oplus\bm{N}^{i}_{v})\bm{W}^{2}),\text{where }\bm{N}^{i}_{v}[j]=\frac{1}{|\mathcal{N}_{j}|}\sum_{k\in N_{j}}\hat{\bm{O}}_{v}^{i}[k]$ (7)
where $\bm{W}^{1}$ and $\bm{W}^{2}$ are weight matrices, $\oplus$ denotes concatenation, $\mathcal{N}_{j}$ represents node entity $j$’s neighbors, $\bm{N}_{v}^{i}$ aggregates neighbor information, $\mathcal{A}_{old}$ is the previous batch’s learned causal graph, and $\Delta\mathcal{A}_{v}\in\mathbb{R}^{n\times n}$ is a learnable adjacency matrix. Unlike $\mathcal{A}_{old}$, $\Delta\mathcal{A}_{v}$ captures unique patterns in the current streaming data batch. $\tilde{\bm{X}}_{v}^{0}$ ($\tilde{\bm{X}}_{v}^{i}$) predicts future values based on previous lagged data $\bm{\hat{X}}_{v}^{0}$ ($\bm{\hat{X}}_{v}^{i}$), leveraging temporal dependencies captured by dilated convolutional neural networks. Finally, we minimize forecasting errors as follows:
$\mathcal{L}_{temporal}$ $=\frac{1}{n(d_{L}+d_{M})}\sum_{v}\sum_{j=1}^{n}\sum_{k=1}^{d_{v}}[||\bm{\hat{X}}_{v}^{0}[j,k]-\tilde{\bm{X}}_{v}^{0}[j,k]||^{2}+||\bm{\hat{X}}_{v}^{i}[j,k]-\tilde{\bm{X}}_{v}^{i}[j,k]||^{2}]$ (8)
Notice that leveraging Eq. 8 and the message passing mechanism of GNNs in Eq. 6 allows the model to encode causality in the learned adjacency matrix $\tilde{\bm{A}}=\mathcal{A}_{old}+\Delta\mathcal{A}_{v}$, such as $\bm{X}\rightarrow\bm{y}$, where $\bm{X}$ is a potential root cause and $\bm{y}$ is a Key Performance Indicator (KPI). Additionally, we add the trace exponential function $h(\tilde{\bm{A}})=(tr(e^{\tilde{\bm{A}}\odot\tilde{\bm{A}}})-n)=0$ as a regularization term to ensure that $\tilde{\bm{A}}$ is acyclic (Pamfil et al., 2020), where $\odot$ denotes the Hadamard product of two matrices.
3.3 Representation Learning with Multi-factor Attention
In microservice systems, each system entity has multiple entity metrics and various log attributes/indicators, including CPU usage, memory usage, log frequency, log golden signal, etc. Existing RCA methods analyze abnormal patterns from each factor (i.e., metric or log indicator) individually, neglecting potential relationships among them. However, the importance of factors varies depending on the abnormal patterns. Hence, reassessing the contribution of each factor to causal structure learning is crucial. To bridge this gap, we propose to explore the correlation of different factors from two modalities and then assess the contribution of each factor to causal structure learning with the attention mechanism (Lu et al., 2016; Vaswani et al., 2017). Given two representations $\bm{H}_{L}^{0}$ and $\bm{H}_{M}^{0}$ in Eq. 3, we compute the multi-factor similarity matrix $\bm{C}_{j}^{0}\in\mathbb{R}^{d_{M}\times d_{L}}$ for historical representation of the $j^{\textrm{th}}$ system entity to capture the correlation of different modalities and the relationship among multiple metrics and log indicators as follows:
$\bm{C}_{j}^{0}$ $=\tanh{(\bm{H}_{M}^{0}[j])^{T}\bm{W}^{3}\bm{H}_{L}^{0}[j])}$ (9)
where $\bm{W}^{3}\in\mathbb{R}^{T_{3}\times T_{3}}$ is a weight matrix and $\bm{H}_{v}^{0}[j]$ denotes the historical representation of the $j^{\textrm{th}}$ system entity for modality $v\in\{M,L\}$. Starting here, we will skip equations related to the $i^{\textrm{th}}$ batch of streaming data for brevity, unless their computation differs from historical data. This matrix measures the similarities between modalities and among multiple factors. By leveraging this similarity matrix, we aim to encode information from both modalities in the hidden representation $\bm{H}_{v}^{0}$ and assess the importance of each factor across both modalities, formulated as:
$\bm{Z}_{L}^{0}[j]$ $=\tanh{(\bm{H}_{L}^{0}[j]\bm{W}^{4}+(\bm{C}_{j}^{0})^{T}\bm{H}_{M}^{0}[j]\bm{W}^{5})}$
$\bm{Z}_{M}^{0}[j]$ $=\tanh{(\bm{H}_{M}^{0}[j]\bm{W}^{5}+\bm{C}_{j}^{0}\bm{H}_{L}^{0}[j]\bm{W}^{4})}$
$\bm{a}_{L}^{0}[j]$ $=\text{softmax}(\bm{w}^{6}\bm{Z}_{L}^{0}[j]),\hskip 8.5359pt\bm{a}_{M}^{0}[j]=\text{softmax}(\bm{w}^{7}\bm{Z}_{M}^{0}[j])$ (10)
@@ 3.3 Representation Learning with Multi-factor Attention
where $\bm{W}^{4}\in\mathbb{R}^{T_{3}\times T_{3}}$ and $\bm{W}^{5}\in\mathbb{R}^{T_{3}\times T_{3}}$ are two weight matrices and $\bm{w}^{6}\in\mathbb{R}^{T_{4}}$ and $\bm{w}^{7}\in\mathbb{R}^{T_{4}}$ are two weight vectors. Notice that $\bm{a}^{0}_{L}[j]$ and $\bm{a}^{0}_{L}[j]$ measure the importance of each factor by encoding information from both modalities, capturing rich relationships for multi-modal and multi-dimensional data. Using these attention vectors, we encode all information learned from multiple factors of two modalities into the weighted representation $\hat{\bm{H}}^{0}_{v}\in\mathbb{R}^{n\times T_{3}}$ by:
$\hat{\bm{H}}^{0}_{v}[j]=\sum_{k=1}^{d_{v}}\bm{a}_{v}^{0}[i,k]\cdot\bm{H}_{v}^{0}[j,k]$ (11)
After encoding the relationship among different factors and two modalities into $\hat{\bm{H}}^{0}_{v}$ and $\hat{\bm{H}}^{i}_{v}$, we aim to recover the factors of two modalities by:
$\bm{O}_{v}^{0}=\text{MLP}^{0}(\hat{\bm{H}}^{0}_{v}),\bm{O}_{v}^{i}=\text{MLP}^{1}(\hat{\bm{H}}^{i}_{v})$ (12)
where $\text{MLP}^{0}$ and $\text{MLP}^{1}$ are two multi-layer perceptrons (MLP) to recover the metrics, $\bm{O}_{v}^{0}\in\mathbb{R}^{n\times d_{v}\times T_{3}}$ and $\bm{O}_{v}^{i}\in\mathbb{R}^{n\times d_{v}\times T_{4}}$. Here, we reshape $\bm{O}_{v}^{0}$ and $\bm{O}_{v}^{i}$ so that $\bm{O}_{v}^{0}\in\mathbb{R}^{n\times d_{v}\cdot T_{3}}$ and $\bm{O}_{v}^{i}\in\mathbb{R}^{n\times d_{v}\cdot T_{4}}$. Overall, Eqs. 9, 10, 11 and 12 are combined to derive $MFL(\cdot)$ in Eq. 5. After we assess the contribution of each factor to the causal structure learning, we use $\bm{a}_{v}[j,k]$ to reweigh the importance of different factors in the future value prediction task in Eq. 8 and further encourage that the representations $\hat{\bm{H}}^{0}_{v}$ and $\hat{\bm{H}}^{i}_{v}$ should contain more information for the factor with a larger weight $\bm{a}_{v}[j,k]$. Therefore, Eq. 8 can be updated as follows:
3.4 Graph Fusion with Contrastive Multi-modal Learning
@@
To tackle the challenges of multi-modal learning (as discussed in challenge C3 in Section 1), we propose to maximize the relatedness between two modalities via contrastive mutual information maximization. Given the representations of historical data $\hat{\bm{H}}^{0}_{v}$ and streaming data $\hat{\bm{H}}^{i}_{v}$ extracted from both metric and log data, we maximize the mutual information between these two modalities:
$\mathcal{L}_{MI}$ $=\mathcal{I}_{\phi}(\hat{\bm{H}}_{M}^{0},\hat{\bm{H}}_{L}^{0})+\mathcal{I}_{\phi}(\hat{\bm{H}}_{M}^{i},\hat{\bm{H}}_{L}^{i})$ (14)
@@
where $\mathcal{I}_{\phi}$ is the mutual information parameterized by a neural network $\phi$. Following InfoNCE style contrastive loss (Oord et al., 2018; Zheng et al., 2022; 2021; 2023; 2024c), we approximate the mutual information with its lower bound as follows:
$\mathcal{I}_{\phi}(\hat{\bm{H}}_{M}^{0},\hat{\bm{H}}_{L}^{0})$ $:=-\frac{1}{n}\sum_{j=1}^{n}\log\frac{\operatorname{sim}(\phi(\hat{\bm{H}}_{M}^{0}[j]),\phi(\hat{\bm{H}}_{L}^{0}[j]))}{\sum_{k}\operatorname{sim}(\phi(\hat{\bm{H}}_{M}^{0}[j]),\phi(\hat{\bm{H}}_{L}^{0}[k]))}$ (15)
@@
where $\operatorname{sim}(a,b)=\exp(\frac{ab^{T}}{|a||b|})$ is the exponential of cosine similarity measurement between two entity representations $a$ and $b$.
To generate the causal graph for the current batch of data, simple addition may not work because it may result in dense and cyclical graphs. This issue might even exacerbate in the low-quality modality scenario, as one modality might convey more important information than others. The low-quality modality usually obscures the crucial patterns for causal graph learning if both modalities are treated with equal importance. To address this issue, we propose to measure the importance of two modalities with the correlation of multiple metrics captured in Eq. 9. Based on the similarity map for the current batch (i.e., $\bm{C}_{j}^{i}$), we further measure the importance of each modality and fuse two causal graphs:
$s_{M}$ $=\frac{\sum_{j=1}^{n}\sum_{l=1}^{d_{M}}\exp(\sum_{k=1}^{d_{L}}\bm{C}^{i}_{j}[l,k])}{\sum_{j=1}^{n}\sum_{l=1}^{d_{M}}\exp(\sum_{k=1}^{d_{L}}\bm{C}^{i}_{j}[l,k])+\sum_{j=1}^{n}\sum_{k=1}^{d_{L}}\exp(\sum_{l=1}^{d_{M}}\bm{C}^{i}_{j}[l,k])}$ (16)
$\mathcal{A}$ $=(1-s_{M})\cdot(\mathcal{A}_{old}+\Delta\mathcal{A}_{L})+s_{M}\cdot(\mathcal{A}_{old}+\Delta\mathcal{A}_{M})$ (17)
Optimization. The final objective function is written as:
$\mathcal{L}_{sparse}$ $=||\Delta\mathcal{A}_{L}||_{1}+||\Delta\mathcal{A}_{M}||_{1}$
$\mathcal{L}$ $=\mathcal{L}_{MI}+\lambda_{1}\mathcal{L}_{temporal}+\lambda_{2}\mathcal{L}_{sparse}+\lambda_{3}h(\mathcal{A})$ (18)
where $||\cdot||_{1}$ is the sparsity constraint imposed on the adjacency matrix and $\mathcal{L}_{sparse}$ aims to ensure that the changes of the edges are expected to be sparse. The trace exponential function $h(\mathcal{A})=(tr(e^{\mathcal{A}\odot\mathcal{A}})-n)=0$ holds if and only if $\mathcal{A}$ is acyclic (Pamfil et al., 2020), where $\odot$ denotes the Hadamard product of two matrices. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the positive constant hyper-parameters.
3.5 Network Propagation-based Root Cause Identification

### v2

@@ III-B Long-term Temporal Causal Relationship Learning
where $\mathbf{f}\in\mathbb{R}^{K}$ represents the 1-D kernel, $\mathbf{f}_{1},\mathbf{f}_{2}\in\mathbb{R}^{K}$ are 1-D convolution kernels, $d$ is the dilation factor, $\odot$ denotes the Hadamard product, $\sigma(\cdot)$ and $\tanh(\cdot)$ are the sigmoid and tanh functions respectively. The output $\mathbf{H}_{v}^{0}\in\mathbb{R}^{n\times d_{v}\times T_{3}}$ represents temporally encoded features for modality $v\in\{M,L\}$. Stacking dilated convolution layers exponentially increases the receptive field, enabling efficient long-term modeling with reduced computational cost. We validate the efficiency of dilated convolutional operations in our experiments (see Subsection IV-B) by comparing their computational costs with those of VAR-based methods.
To explore the correlation of different factors from two modalities and then assess the contribution of each factor to causal structure learning, we compute a multi-factor similarity matrix for the $j^{\textrm{th}}$ system entity:
$\mathbf{C}_{j}^{0}=\tanh(\mathbf{H}_{M}^{0}[j]\mathbf{W}^{3}(\mathbf{H}_{L}^{0}[j])^{\top}),$ (4)
where $\mathbf{W}^{3}\in\mathbb{R}^{T_{3}\times T_{3}}$ is a learnable projection. The similarity matrix $\mathbf{C}_{j}^{0}\in\mathbb{R}^{d_{L}\times d_{M}}$ captures inter-factor relationships between metric and log modalities for the historical data. We then compute attention-based importance weights to quantify each factor’s contribution to causal structure learning:
$\mathbf{Z}_{L}^{0}[j]$ $=\tanh(\mathbf{H}_{L}^{0}[j]\mathbf{W}^{4}+\mathbf{H}_{M}^{0}[j]\mathbf{W}^{5}\mathbf{C}_{j}^{0}),$
$\mathbf{Z}_{M}^{0}[j]$ $=\tanh(\mathbf{H}_{M}^{0}[j]\mathbf{W}^{5}+\mathbf{H}_{L}^{0}[j]\mathbf{W}^{4}(\mathbf{C}_{j}^{0})^{\top}),$ (5)
$\mathbf{a}_{L}^{0}[j]$ $=\text{softmax}(\mathbf{w}^{6}\mathbf{Z}_{L}^{0}[j]),\quad\mathbf{a}_{M}^{0}[j]=\text{softmax}(\mathbf{w}^{7}\mathbf{Z}_{M}^{0}[j]),$
@@
where $\mathbf{W}^{4},\mathbf{W}^{5}\in\mathbb{R}^{T_{3}\times T_{3}}$ and $\mathbf{w}^{6},\mathbf{w}^{7}\in\mathbb{R}^{T_{4}}$. The attention vectors $\mathbf{a}_{v}^{0}[j]$ measure the importance of each factor by encoding information from both modalities, capturing rich relationships for multi-modal and multi-dimensional data. With these attention vectors, the weighted modality representations are obtained as:
$\hat{\mathbf{H}}_{v}^{0}[j]=\sum_{k=1}^{d_{v}}\mathbf{a}_{v}^{0}[j,k]\cdot\mathbf{H}_{v}^{0}[j,k],$ (6)
and then passed through a multi-layer perceptron (MLP) for factor recovery:
$\mathbf{O}_{v}^{0}=\text{MLP}^{0}(\hat{\mathbf{H}}_{v}^{0}).$ (7)
Overall, the multi-factor learning process (Eqs. (4), (5), (6) and (7)) is summarized as:
$\hat{\mathbf{O}}_{v}^{0}=MFL(\hat{\mathbf{H}}_{v}^{0}).$ (8)
To learn causal relationships among entities, we apply a message-passing GNN (GraphSAGE [hamilton2017inductive]) to mimic fault propagation through a message-passing mechanism:
$\tilde{\mathbf{X}}_{v}^{0}$ $=\sigma_{2}(\mathcal{A}_{\text{old}}(\hat{\mathbf{O}}_{v}^{0}\oplus\mathbf{N}_{v}^{0})\mathbf{W}^{1}),$ (9)
$\mathbf{N}_{v}^{0}[j]$ $=\frac{1}{|\mathcal{N}_{j}|}\sum_{k\in\mathcal{N}_{j}}\hat{\mathbf{O}}_{v}^{0}[k],$
where $\mathbf{W}^{1}$ is the weight matrix, $\sigma_{2}$ denotes ReLU, $\oplus$ denotes concatenation, $\mathcal{A}_{\text{old}}$ is the historical causal graph, $\mathcal{N}_{j}$ represents node entity $j$’s neighbors, and $\mathbf{N}_{v}^{0}$ aggregates neighbor information. $\tilde{\mathbf{X}}_{v}^{0}$ predicts future values based on previous lagged data $\mathbf{\hat{X}}_{v}^{0}$, leveraging temporal dependencies captured by dilated convolutional neural networks. Finally, the training objective minimizes the forecasting error:
$\mathcal{L}_{t}=\frac{1}{n(d_{L}+d_{M})}\sum_{v}\sum_{j=1}^{n}\sum_{k=1}^{d_{v}}\mathbf{a}_{v}^{0}[j,k]\|\hat{\mathbf{X}}_{v}^{0}[j,k]-\tilde{\mathbf{X}}_{v}^{0}[j,k]\|^{2}.$ (10)
The $\mathbf{a}_{v}^{0}[j,k]$ emphasize influential factors during temporal forecasting, reinforcing interpretable causal learning.
Dynamic Encoder.
The Dynamic Encoder extends the same architecture to model streaming data in the online setting, enabling adaptive causal updates. It aims to efficiently model the long-term temporal dependencies and the causal relations among system entities and KPIs for the streaming data in the online setting. Given the current batch data $\hat{\mathbf{X}}_{v}^{i}$, temporal dependencies are encoded as:
$\mathbf{H}_{v}^{i}$ $=\tanh(g(\hat{\mathbf{X}}_{v}^{i},\mathbf{f}_{3}))\odot\sigma(g(\hat{\mathbf{X}}_{v}^{i},\mathbf{f}_{4})),$ (11)
$\hat{\mathbf{O}}_{v}^{i}$ $=MFL(\hat{\mathbf{H}}_{v}^{i}),$ (12)
where $\mathbf{f}_{3},\mathbf{f}_{4}$ are dilated convolutional kernels. We then propagate messages through the updated causal graph:
$\tilde{\mathbf{X}}_{v}^{i}$ $=\sigma_{2}((\mathcal{A}_{\text{old}}+\Delta\mathcal{A}_{v})(\hat{\mathbf{O}}_{v}^{i}\oplus\mathbf{N}_{v}^{i})\mathbf{W}^{2}),$ (13)
$\mathbf{N}_{v}^{i}[j]$ $=\frac{1}{|\mathcal{N}_{j}|}\sum_{k\in\mathcal{N}_{j}}\hat{\mathbf{O}}_{v}^{i}[k],$ (14)
where $\Delta\mathcal{A}_{v}\in\mathbb{R}^{n\times n}$ is added to capture incremental causal changes in the current batch. Finally, we jointly minimize the reconstruction losses from historical and streaming data:
$\mathcal{L}_{t}$ $=\frac{1}{n(d_{L}+d_{M})}\sum_{v}\sum_{j=1}^{n}\sum_{k=1}^{d_{v}}\mathbf{a}_{v}^{i}[j,k]\|\hat{\mathbf{X}}_{v}^{i}[j,k]-\tilde{\mathbf{X}}_{v}^{i}[j,k]\|^{2}$
By integrating Eq. (III-B) with the message-passing mechanism in Eq. (13), the model continuously refines the causal adjacency matrix $\tilde{\mathcal{A}}=\mathcal{A}_{\text{old}}+\Delta\mathcal{A}_{v}$, capturing dynamic causal relations such as $\mathbf{X}\rightarrow\mathbf{y}$, where $\mathbf{X}$ denotes a potential root cause and $\mathbf{y}$ is the system KPI.
III-C Graph Fusion with Contrastive Multi-modal Learning
@@
To address the challenges of multi-modal learning (as discussed in Challenge C3 in Section I), we propose to enhance the relatedness between two modalities through contrastive mutual information maximization. Given the representations of historical data $\hat{\mathbf{H}}^{0}_{v}$ and current batch data $\hat{\mathbf{H}}^{i}_{v}$ extracted from both metric and log data, we maximize the mutual information between these modalities:
$\mathcal{L}_{\mathrm{MI}}$ $=\mathcal{I}_{\phi}(\hat{\mathbf{H}}_{M}^{0},\hat{\mathbf{H}}_{L}^{0})+\mathcal{I}_{\phi}(\hat{\mathbf{H}}_{M}^{i},\hat{\mathbf{H}}_{L}^{i}),$ (15)
@@
where $\mathcal{I}_{\phi}$ is the mutual information parameterized by a neural network $\phi$. Following the InfoNCE-style contrastive loss [oord2018representation, DBLP:conf/www/ZhengBWZH25, DBLP:journals/tmlr/ZhengFMH24, DBLP:conf/kdd/ZhengXZH22, DBLP:journals/corr/abs-2411-15623, DBLP:conf/sdm/ZhengZH23], we approximate the mutual information by its lower bound:
$\mathcal{I}_{\phi}(\hat{\mathbf{H}}_{M}^{0},\hat{\mathbf{H}}_{L}^{0})$ $:=\frac{1}{n}\sum_{j=1}^{n}\log\frac{\mathrm{sim}(\phi(\hat{\mathbf{H}}_{M}^{0}[j]),\phi(\hat{\mathbf{H}}_{L}^{0}[j]))}{\sum_{k}\mathrm{sim}(\phi(\hat{\mathbf{H}}_{M}^{0}[j]),\phi(\hat{\mathbf{H}}_{L}^{0}[k]))},$ (16)
@@
where $\mathrm{sim}(a,b)=\exp\!\left(\frac{ab^{\top}}{\|a\|\|b\|}\right)$ represents the exponential of cosine similarity between two entity representations $a$ and $b$.
To generate the causal graph for the current batch, directly summing the graphs from both modalities may yield overly dense or cyclic structures. This problem is amplified when one modality is of lower quality, as treating both modalities equally can obscure critical causal patterns. To mitigate this issue, we estimate the relative importance of the two modalities based on the correlation of multiple metrics. Using the similarity map for the current batch (i.e., $\mathbf{C}_{j}^{i}$), we compute the modality importance and fuse the two causal graphs as follows:
$\mathbf{C}_{j}^{i}$ $=\tanh\!\big(\mathbf{H}_{M}^{i}[j]\mathbf{W}^{8}(\mathbf{H}_{L}^{i}[j])^{\top}\big),$
$s_{M}$ $=\frac{\sum_{j}\sum_{l}\exp\!\big(\sum_{k}\mathbf{C}_{j}^{i}[l,k]\big)}{\sum_{j}\Big[\sum_{l}\exp\!\big(\sum_{k}\mathbf{C}_{j}^{i}[l,k]\big)+\sum_{k}\exp\!\big(\sum_{l}\mathbf{C}_{j}^{i}[l,k]\big)\Big]},$
$\mathcal{A}$ $=(1-s_{M})\big(\mathcal{A}_{\mathrm{old}}+\Delta\mathcal{A}_{L}\big)+s_{M}\big(\mathcal{A}_{\mathrm{old}}+\Delta\mathcal{A}_{M}\big),$
where $\mathbf{W}^{8}$ is a learnable weight matrix. This adaptive fusion mechanism ensures that higher-quality modalities contribute more to the causal graph update while maintaining structural sparsity and stability.
Optimization. The overall training objective is defined as:
$\mathcal{L}_{\text{sparse}}$ $=\|\Delta\mathcal{A}_{L}\|_{1}+\|\Delta\mathcal{A}_{M}\|_{1},$
$\mathcal{L}$ $=-\mathcal{L}_{MI}+\lambda_{1}\mathcal{L}_{t}+\lambda_{2}\mathcal{L}_{\text{sparse}}+\lambda_{3}h(\mathcal{A}),$ (17)
where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm that enforces sparsity on the changes of the adjacency matrices. The sparsity regularization $\mathcal{L}_{\text{sparse}}$ encourages only a limited number of edge updates across modalities. The acyclicity constraint $h(\mathcal{A})=\mathrm{tr}(e^{\mathcal{A}\odot\mathcal{A}})-n=0$ holds if and only if $\mathcal{A}$ is a directed acyclic graph [DBLP:conf/aistats/PamfilSDPGBA20], where $\odot$ denotes the Hadamard product. $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are positive hyperparameters balancing the contribution of each component.
Online Finetuning. During the offline training stage, both the static and dynamic encoders are jointly optimized using Eq. (III-C). In the online adaptation stage, the static encoder is frozen while the dynamic encoder is finetuned to efficiently adapt to streaming data with reduced computational overhead.
III-D Network Propagation-based Root Cause Identification